What discount rate would make you indifferent between receiving $3,393.00 per year forever and $5,267.00 per year for 23.00 years? Assume the first payment of both cash flow streams occurs in one year.
Answer #1

PV of Cash flow forever = PV of Annuity for 23 years
3393/r = 5267*(1-(1+r)-23)/r
3393 = 5267 (1-(1+r)-23
(1+r)-23 = 0.3558
1+r = (1/0.3558)1/23
r = 4.595% or 4.60%

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well